Hello, can you please answer it and show me step by step how to do it? Than you.
Fred mixed 9 ml of a salt solution with 12 ml of a 47% salt solution, to make a 41% salt solution. Find the percent salt concentration of the first solution.

Answer by stanbon(75887)

Fred mixed 9 ml of a salt solution with 12 ml of a 47% salt solution, to make a 41% salt solution. Find the percent salt concentration of the first solution.
------------
Equation:
salt + salt = salt
(x/100)*9 + 0.47*12 = 0.41(9+12)
------
9x = 41*9 +41*12 - 47*12
9x = 41*9 - 6*12
9x = 297
x = 33% (percent salt in the first solution)
---------
Cheers,
Stan H.